OpenStudy (anonymous):

A quadrilateral has three angles that measure 80°, 110°, and 75°. Which is the measure of the fourth angle?Select one of the options below as your answer: A. 50° B. 90° C. 95° D. 125°

OpenStudy (anonymous):

okay well how many degrees are in any quadrilateral? :)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

they should all add up to 360 degrees so therefore you make an equation let x be the missing angle measure 80+110+75+x=360 combine the like terms 265+x=360 move the 265 over the right side x=??

OpenStudy (anonymous):

do you get what im doing here? :)

OpenStudy (anonymous):

@AshleyGriselle26

OpenStudy (anonymous):

x+265=360 x=360-265 x=95

OpenStudy (anonymous):

o ok i understand nw thanks

OpenStudy (anonymous):

work it and tell me what you get :)

OpenStudy (anonymous):

thanks i got 95

OpenStudy (anonymous):

yay! :D

OpenStudy (anonymous):

well i hope i helped...message me if you have anymore questions! :)

OpenStudy (anonymous):

anything else i can help with? (:

OpenStudy (anonymous):

:DD Thanks and no problem

OpenStudy (anonymous):

k good luck! :D

OpenStudy (anonymous):

thank u :DD

OpenStudy (anonymous):

anytime! :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

:)